4./10 points SerPSE10 3.4.0P012 MI Consider the two vectors A-5 i-j and B-i-2 j. (a) Calculate A +B (b) Calculate A B (c) Calculate A-B 1 (d) Calculate A B (e) Calculate the directions of A+ B and A -B. o (counterclockwise from the +x axis) o(counterclockwise from the +x axis) A + B A -B
Answer #1

Solution)

Given, A= 5i-j

B=-i-2j

A) A+B= (5i-j) +(-i-2j) =4i-3j (Ans)

======

B) A-B = (5i-j)-(-i-2j) = 6i+j (Ans)

=====

C) mod(A+B)=sqrt(4^2 + 3^2) =5 (Ans)

=====

D) mod(A-B) = sqrt(6^2 + 1^2) = 6.08 (Ans)

=======